OpenStudy (arbyj):

how do i solve 1/2(x-4)^2=8 by square rooting?

OpenStudy (mathmale):

Do you mean (1/2)(x-4)^2=8? or do you mean 1 / [2(x-4)^2=8? which?

OpenStudy (arbyj):

the first one

OpenStudy (mathmale):

Thank you. Please multiply both sides of this equation by 2 (for the purpose of getting rid of the fraction 1/2):\[\frac{ 1 }{ 2 }(x-4)^2=8\]

OpenStudy (mathmale):

\[\frac{ 1 }{ 2 }(x-4)^2=8\]

OpenStudy (mathmale):

Your result?

OpenStudy (arbyj):

i already did those steps but im stuck at \[(x-4)^2 = \sqrt{16}\]

OpenStudy (mathmale):

Where did the square root over the 16 come from? If you find the sqrt of the right side (16) of your equation, you must also find the sqrt of the left side. Please show this as an equation.

OpenStudy (arbyj):

i was doing it the way i was shown

OpenStudy (arbyj):

after doing that, i got \[x -4 =\pm 4\]

OpenStudy (arbyj):

is that correct?

OpenStudy (mathmale):

sorry, but your \[(x-4)^2 = \sqrt{16}\]

OpenStudy (mathmale):

is not correct. If you find the sqrt of 16, you must also find the sqrt of (x-4)^2 at the same time (not later on).

OpenStudy (mathmale):

what is \[\sqrt{(x-4)^2}?\]

OpenStudy (mathmale):

What is \[\sqrt{16}?\]

OpenStudy (arbyj):

4

OpenStudy (mathmale):

The set of answers comes from \[\sqrt{(x-4)^2}=\pm \sqrt{16}\]

OpenStudy (mathmale):

Simplify the left side and set the new left side equal to \[\pm \sqrt{16}\]

OpenStudy (arbyj):

i cant see the equations anymore

OpenStudy (mathmale):

Simplify the following:\[\sqrt{(x-4)^2}=\pm \sqrt{16}\]

OpenStudy (mathmale):

\[\sqrt{(x-4)^2}=\pm \sqrt{16}\]

OpenStudy (mathmale):

What does the left side simplify to?

OpenStudy (arbyj):

but i calculated and still got this 

OpenStudy (mathmale):

Yes, and on the right size you have \[\pm4\]

OpenStudy (mathmale):

So, x-4 = plus or minus 4 Solve that for x, please (2 answers required)

OpenStudy (arbyj):

so i was correct the first time but i was just missing the part where i square root (x-4)^2?

OpenStudy (mathmale):

Looks that way. Please, would you finish the problem solution in the manner I've outlined above?